OpenStudy (anonymous):

5x + 10y = 18 x = y - 6 Which of the following equations is the result of substituting y - 6 in for x in the first equation?

OpenStudy (anonymous):

Instead of an 'x', put in 'y-6' into the first equation. Use brackets to stay organized. 5(y-6) + 10y = 18 Then you can expand it.

OpenStudy (texaschic101):

just distribute the 5 through the parenthesis in jam's response to get the result

OpenStudy (anonymous):

a) 5x + 10y - 6 = 18 b) 5y - 6 + 10y = 18 c) 5y - 30 + 10y = 18

OpenStudy (texaschic101):

5(y - 6) + 10y = 18... this is the same as : 5 * y - 5 * 6 + 10y = 18 can you do it now ?

OpenStudy (texaschic101):

ok...I am feeling generous...answer is 5y - 30 + 10y = 18

OpenStudy (anonymous):

thank you. Will you help me with my others?